OpenStudy (anonymous):

what is halfway between 123 and 321 please include a complete explanation and if u want a drawing :(

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

321 - 123 = 198 198/2 =99 123 + 99 =222 Answer